OpenStudy (anonymous):

Can you have a right triangle where the lengths of the legs are whole numbers and the length of the hypotenuse is 23? Explain.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@amistre64

OpenStudy (inkyvoyd):

this is solving the diophantine equation for a^2+b^2=c^2 where c=23

OpenStudy (anonymous):

yes i am given a = x^2-y^2 and b = 2xy for generating pythagorean triples

OpenStudy (anonymous):

this problem is from polynomial identities